# A Survey of Privacy-Preserving Model Explanations: Privacy Risks, Attacks, and Countermeasures

### v1

@@ 1. Introduction
Figure 1. This work vs. existing surveys. Explainable AI involves explanation and interpretable methods (e.g. (Bodria et al., 2023; Guidotti et al., 2018; Gilpin et al., 2018)). Adversarial AI includes adversarial attacks on ML models (e.g. (Machado et al., 2021; Biggio and Roli, 2018)). Privacy AI involves privacy issues in ML (e.g. (Rigaki and Garcia, 2023; Hu et al., 2022; Liu et al., 2021)). Others (Ferry et al., 2023b; Baniecki and Biecek, 2024) discuss exploits on model explanations. Our survey offers the first complete picture on privacy attacks, leaks, and defenses in explainable AI.
@@ 1.1. Comparisons with existing surveys
Many surveys have summarised different privacy issues on ML models (Biggio and Roli, 2018; Papernot et al., 2017; Machado et al., 2021; Liu et al., 2021), while others reviewed explanation methods for ML models (Gilpin et al., 2018; Bodria et al., 2023; Adadi and Berrada, 2018), but not both. For example, Rigaki et al. (Rigaki and Garcia, 2023) presented a thorough analysis of over 45 publications on privacy attacks in machine learning, spanning the last seven years. Hu et al. (Hu et al., 2022) surveyed a special type of privacy attacks, called membership inference. On the other hand, others (Guidotti et al., 2018; Adadi and Berrada, 2018; Došilović et al., 2018) offered a comprehensive classification of model explanations to enhance interpretability and provided guidance for selecting suitable methods for specific ML models and desired explanations.
Some existing surveys summarised adversarial attacks but presented partial coverage of privacy attacks on model explanations with basic introductions and limited discussions of the methods. Ferry et al. (Ferry et al., 2023b) examined the interplay between interpretability, fairness, and privacy, which are critical for responsible AI, particularly in high-stakes decision-making like college admissions and credit scoring. Baniecki et al. (Baniecki and Biecek, 2024) surveyed adversarial attacks on model explanations and fairness metrics, offered a unified taxonomy for clarity across related research areas, and discussed defensive strategies against such attacks. However, these papers are either too high level or too specialised in non-privacy attacks.
@@ 1.2. Paper collection methodology
Finding relevant research on this subject proved to be complex due to its incorporation of various topics such as data privacy, privacy attacks, explanations of models, explainable AI (XAI), and the development of privacy-preserving explanations. To navigate this breadth of concepts, we employed diverse keyword combinations pertaining to “privacy”, “explanation”, and specific attack types including “membership inference”, “data reconstruction”, “attribute inference”, “model extraction”, “model stealing”, “property inference”, and “model inversion”. Our initial search utilised platforms like Google Scholar, Semantic Scholar, and Scite.ai – an AI-enhanced search tool – to assemble a preliminary collection of studies. This selection was expanded through backward searches, analysing the references of initially chosen papers, and forward searches, identifying papers that cited the initial ones. Additionally, we manually verified the relevance and focus of these articles across various sources due to discrepancies, such as some studies addressing privacy in the context of safeguarding against manipulation attacks instead of privacy intrusions. Ultimately, this process culminated in identifying nearly 50 pivotal research papers on the topic.
@@ 1.3. Contributions of the article
  • Connected Taxonomies: We have organised all existing literature on PrivEx according to various criteria, including the types of explanations targeted and the methodologies employed in attacks and defenses. Fig. 2 showcases the taxonomy we have developed to structure these works.
@@
  • Challenges and Future Directions: Designing privacy-preserving explanations for machine learning models is an emerging field of research. From the surveyed literature, we highlight unresolved issues and suggest a number of potential research directions into both the offensive and defensive aspects of privacy in model explanations.
@@ 1.4. Organisation of the article
The rest of the article is organised as follows. § 2 revisits model explanations, acting as foundations for privacy attacks. § 3 presents the taxonomy of privacy attacks on model explanations and provides in-depth descriptions, including threat model and attack scenarios. § 4 discusses the causes of privacy leaks in model explanations. § 5 explores countermeasures and a new class of privacy-preserving model explanations by design. § 6 provides the pinpoints to existing resources including source code, datasets, and evaluation metrics. Finally, § 7 contains a discuss on ongoing and upcoming research directions and § 8 concludes the survey.
@@ 2.1. Feature-based Explanations
Figure 4. Comparison of the four feature-based explanation approaches: vanilla gradients, integrated gradients, guided backpropagation, and SmoothGrad. SmoothGrad shows better discriminativity than the other methods. The figure is taken from  (Smilkov et al., 2017).
Perturbation-based. Perturbation-based explanations involve querying a model that needs to be explained with a series of altered inputs (Shokri et al., 2021). SmoothGrad (Smilkov et al., 2017) is a popular perturbation-based explanation method that produces several samples by injecting Gaussian noise into the input data and then computes the mean of the gradients from these samples (See Fig. 4). Formally, for a certain $k$ samples, the explanation function is defined as:
@@ 2.2. Interpretable Surrogates
Figure 5. Examples of explaining a learnt model using LIME. The figure is taken from (Ribeiro et al., 2016).
@@
where $X^{\prime}$ is the neighbourhood of $x$. For example, as shown in Fig. 5, LIME can provide pixel-wise explanations to the model predictions based on a given image.
@@
for each feature index $i=1,2,\ldots,M$, where $x$ is drawn from the foreground $F$, and $x^{0}$ from the background $B$.
@@ 2.3. Example-based Explanations
Figure 6. Examples of learnt prototypes from the MNIST image dataset. The figure is taken from (Li et al., 2018).
Intrinsic methods for deep learning. In deep learning, intrinsic interpretability can be provided by prototype-based or distance-based methods (Montenegro et al., 2022). For instance, the Explainable Deep Neural Network (xDNN) (Angelov and Soares, 2020b) and Deep Machine Reasoning (DMR) (Angelov and Soares, 2020a) define prototypes as dense data points and classify observations based on the closest prototype. The Prototype Classifier method learns representative prototypes from training data, using an autoencoder for feature extraction and classification based on latent representations (Li et al., 2018) (see Fig. 6). The Prototypical Part Network (ProtoPNet) represents image parts in clusters in a latent space, which are used to predict and explain classifications (Chen et al., 2019). Additionally, the Deep k-Nearest Neighbors (DkNN) calculates neighbors at each model layer to ensure consistent predictions, offering explanations based on similar examples across the model’s entirety (Papernot and McDaniel, 2018).
@@ 2.4. Counterfactual Explanations
Counterfactual explanations (aka algorithmic recourse) provide insights into how slight changes to input features could lead to different model outcomes, aiding in tasks like model debugging and ensuring regulatory compliance (Goethals et al., 2023; Kuppa and Le-Khac, 2021). The study in (Kuppa and Le-Khac, 2021) gives an illustration of counterfactual and other four sample categories (i. e., adversarial examples, local robustness, invariant samples, and uncertainty samples) through the boundaries between human analysi and a learnt model. The application of counterfactual explanations varies with the model’s complexity and includes considerations such as model transparency, type compatibility, and adherence to constraints like feasibility and causality (Wachter et al., 2017; Dodge et al., 2019; Binns et al., 2018). The concept overlaps with other areas of research such as algorithmic recourse, inverse classification, and contrastive explanations (Karimi et al., 2021; Ustun et al., 2019; Laugel et al., 2017; Dhurandhar et al., 2018).
Figure 7. Decision boundaries between human analyst and a learnt model. The figure is taken from (Kuppa and Le-Khac, 2021).
@@
Diverse counterfactuals. Recent works study the generation of multiple alternative counterfactuals per input, offering a spectrum of potential changes rather than just one nearest option (Mothilal et al., 2020). This approach empowers users by offering them various ways they could potentially modify their data to achieve a preferred result.
@@ 3.1. Membership Inference Attacks (MIA)
Threat model. The adversary is able to submit $x$ to the black-box model (Liu et al., 2022d; Li et al., 2022; Carlini et al., 2022; Ye et al., 2022) to receive the prediction $f(x)$ and any corresponding explanations, despite not having direct access to the model’s internals (Quan et al., 2022) (see Fig. 8). However, they are assumed to know the model’s architecture and possess an auxiliary dataset similar to the model’s training data, reflected in much of the current research on the topic (Liu et al., 2024b).
@@
Figure 9. Model-based membership inference attacks proposed in (Liu et al., 2024b).
Liu et al. (Liu et al., 2024b) proposes a model-based attack that involves four main stages: training a shadow model, extracting attribution features, training an attack model, and inferring membership (see Fig. 9). The adversary starts by training a shadow model using an auxiliary dataset that is similar to the training data of the target model. Then, attribution maps are generated for a given sample, and perturbations are applied based on these maps to observe changes in predictions. Next, the adversary trains an attack model, typically a Multi-Layer Perceptron (MLP), using the attribution features combined with other data such as loss values and one-hot encoded class information to construct features indicative of membership.
@@ 3.3. Reconstruction Attacks
  • Threat model: We consider a machine learning model $f_{t}$ that processes confidential data $x$ from a set $X_{p}$ (for instance, facial images). It employs these private inputs to generate a prediction $\hat{y}_{t}$ (such as identifying emotions). An issue arises when an attacker gains access to the target prediction $\hat{y}_{t}$ and the explanation $\phi_{t}$ (due to reasons like a data breach, interception during transmission, or sharing on social media). One scenario is to assume that the attacker only has the compromised data, an independent dataset $x\in X_{a}$, and the ability to interact with the target model via black-box (Zhao et al., 2021b). The attacker does not require additional privileged information, such as blurred versions of the images. The objective of the attacker is to develop their own inversion model $f_{a}$ to reconstruct the original image $x$ from the model’s outputs $\hat{y}_{t},\phi_{t}$). Such a reconstruction would allow them to predict sensitive information from the reconstructed image $\hat{x}_{r}$, including the possibility of re-identifying the individual from the facial emotion recognition system.
@@ 3.5. Model Extraction Attacks
Research on model extraction attacks targeting explainable AI systems is emerging. Milli et al. (Milli et al., 2019) develops a method that leverages the discrepancy in gradient-based explanations between an original AI model and its clone, demonstrating enhanced attack efficiency. Additionally, Ulrich et al. (Aïvodji et al., 2020) designs an attack utilising counterfactual explanations to train a cloned model with greater effectiveness. Miura et al. (Miura et al., 2021) designs a data-free attack that does not require surrogate datasets in advance.
@@ 4.1. Privacy Leaks in Counterfactual Explanations
Diverse counterfactuals equip users with a range of actionable insights to potentially alter their outcomes favorably (Mothilal et al., 2020; Nguyen et al., 2023a; Thang et al., 2015; Nguyen et al., 2015a). However, this also increases privacy risks as it may give away additional details that could be exploited for more potent attacks (Aïvodji et al., 2020). Artelt et al. (Artelt et al., 2021) identifies a key problem with counterfactual explanations: their instability to minor input variations can lead to significantly different outcomes for similar cases. Addressing this, the authors propose studying the robustness of counterfactual explanations and suggest using plausible rather than closest counterfactuals to enhance stability (Artelt and Hammer, 2020).
@@ 4.4. Causes of Property Inference Attacks
Regularisation techniques like dropout and ensemble learning have been shown to prevent models from memorizing private inputs, potentially reducing the risk of information leakage (Luo et al., 2021; Melis et al., 2019). Despite previous findings, Luo et al. (Luo et al., 2022) reveals that incorporating dropout in neural networks at varying rates (0.2, 0.5, 0.8) actually enhances the accuracy of certain attacks. This counterintuitive result is attributed to dropout preventing overfitting by smoothing the decision boundaries, which inadvertently benefits the attack. Nevertheless, a very high dropout rate (0.8) does decrease the success rates of one attack due to underfitting and increased randomness in the model, which disrupts the linearity between inputs and outputs.
@@ 5.1. Defences with Differential Privacy
Differential privacy (DP) is a solid, mathematically based privacy standard that defines privacy loss using a quantifiable metric. It does so through mechanisms that guarantee the aggregated data output will obscure the involvement of any individual record in the dataset, as established by Dwork et al. (Dwork et al., 2014). Differential privacy is usually formalized as follows (Huang et al., 2023). A randomized mechanism $M$ with domain $D$ and range $R$ achieves $\varepsilon$-differential privacy ($\varepsilon$-DP) if, for all adjacent datasets $d,d^{\prime}$ differing by one row, and for any output set $S\subseteq R$, the following inequality holds:
@@ 5.1.3. DP-Locally Linear Maps
To create differentially private Locally Linear Maps (LLM), Harder et al. (Harder et al., 2020) employs the moments accountant technique combined with differentially private stochastic gradient descent (DP-SGD) (Abadi et al., 2016). The perturbation process involves two main steps per iteration for each minibatch of size $L$: (1) Clipping the norm of the datapoint-wise gradient $h_{t}(x_{n})$ using a threshold $C$ and adding Gaussian noise to it, resulting in $\hat{h}_{t}$: $\hat{h}_{t}\leftarrow\frac{1}{L}\sum_{n=1}^{L}h_{t}(x_{n})+\mathcal{N}(0,\sigma^{2}C^{2}I)$. (2) Updating the LLM parameters in the descending direction: $W_{t+1}\leftarrow W_{t}-\eta\hat{h}_{t}$. This process ensures that the final LLM is $(\epsilon,\delta)$-differentially private. To improve the privacy-accuracy trade-off, especially for high-dimensional inputs like images, the author suggest reducing the dimensionality of the parameters by first projecting them onto a lower-dimensional space using a shared matrix $R_{m}$, and then perturbing the gradients of the projected parameters.

### v2

@@ 1. Introduction
Figure 1. This work vs. existing surveys. Explainable AI involves explanation and interpretable methods (e.g. (Bodria et al., 2023; Guidotti et al., 2018; Gilpin et al., 2018)). Adversarial AI includes adversarial attacks on ML models (e.g. (Machado et al., 2021; Biggio and Roli, 2018)). Privacy AI involves privacy issues in ML (e.g. (Rigaki and Garcia, 2023; Hu et al., 2022b; Liu et al., 2021)). Others (Ferry et al., 2023b; Baniecki and Biecek, 2024) discuss exploits on model explanations. Our survey offers the first complete picture on privacy attacks, leaks, and defenses in explainable AI.
@@ 1.1. Comparisons with existing surveys
Many surveys have summarised different privacy issues on ML models (Biggio and Roli, 2018; Papernot et al., 2017; Machado et al., 2021; Liu et al., 2021), while others reviewed explanation methods for ML models (Gilpin et al., 2018; Bodria et al., 2023; Adadi and Berrada, 2018), but not both. For example, Rigaki et al. (Rigaki and Garcia, 2023) presented a thorough analysis of over 45 publications on privacy attacks in machine learning, spanning the last seven years. Hu et al. (Hu et al., 2022b) surveyed a special type of privacy attacks, called membership inference. On the other hand, others (Guidotti et al., 2018; Adadi and Berrada, 2018; Došilović et al., 2018) offered a comprehensive classification of model explanations to enhance interpretability and guided the selection of suitable methods for specific ML models and desired explanations.
Some existing surveys summarised adversarial attacks but presented partial coverage of privacy attacks on model explanations with basic introductions and limited discussions of the methods. Ferry et al. (Ferry et al., 2023b) examined the interplay between interpretability, fairness, and privacy, which are critical for responsible AI, particularly in high-stakes decision-making like college admissions and credit scoring. Baniecki et al. (Baniecki and Biecek, 2024) surveyed adversarial attacks on model explanations and fairness metrics, offered a unified taxonomy for clarity across related research areas, and discussed defensive strategies against such attacks. However, these papers are either too high-level or too specialised in non-privacy attacks.
@@ 1.2. Paper collection methodology
Finding relevant research on this subject proved to be complex due to its incorporation of various topics such as data privacy, privacy attacks, explanations of models, explainable AI (XAI), and the development of privacy-preserving explanations. To navigate this breadth of concepts, we employed diverse keyword combinations about “privacy”, “explanation”, and specific attack types including “membership inference”, “data reconstruction”, “attribute inference”, “model extraction”, “model stealing”, “property inference”, and “model inversion”. Our initial search utilised platforms like Google Scholar, Semantic Scholar, and Scite.ai – an AI-enhanced search tool – to assemble a preliminary collection of studies. This selection was expanded through backward searches, analysing the references of initially chosen papers, and forward searches, identifying papers that cited the initial ones. Additionally, we manually verified the relevance and focus of these articles across various sources due to discrepancies, such as some studies addressing privacy in the context of safeguarding against manipulation attacks instead of privacy intrusions. Ultimately, this process culminated in nearly 50 pivotal research papers on the topic.
@@ 1.3. Contributions of the article
  • Connected Taxonomies: We have organised all existing literature on PrivEx according to various criteria, including the types of explanations targeted and the methodologies employed in attacks and defences. Fig. 2 showcases the taxonomy we have developed to structure these works.
@@
  • Challenges and Future Directions: Designing privacy-preserving explanations for machine learning models is an emerging field of research. From the surveyed literature, we highlight unresolved issues and suggest several potential research directions into both the offensive and defensive aspects of privacy in model explanations.
@@ 1.4. Organisation of the article
The rest of the article is organised as follows. § 2 revisits model explanations, acting as foundations for privacy attacks. § 3 presents the taxonomy of privacy attacks on model explanations and provides in-depth descriptions, including threat model and attack scenarios. § 4 discusses the causes of privacy leaks in model explanations. § 5 explores countermeasures and a new class of privacy-preserving model explanations by design. § 6 provides the pinpoints to existing resources including source code, datasets, and evaluation metrics. Finally, § 7 contains a discussion on ongoing and upcoming research directions and § 8 concludes the survey.
@@ 2.1. Feature-based Explanations
Perturbation-based. Perturbation-based explanations involve querying a model that needs to be explained with a series of altered inputs (Shokri et al., 2021). SmoothGrad (Smilkov et al., 2017) is a popular perturbation-based explanation method that produces several samples by injecting Gaussian noise into the input data and then computes the mean of the gradients from these samples.Formally, for a certain $k$ samples, the explanation function is defined as:
@@ 2.2. Interpretable Surrogates
where $X^{\prime}$ is the neighbourhood of $x$.
@@
for each feature index $i=1,2,\ldots,M$. In other words, to conduct a global analysis of model behavior, it is necessary to consider predictions at multiple inputs $x\sim\mathcal{F}$ from a distribution $\mathcal{F}$ called the foreground. Since the choice of baseline $x^{0}$ is ambiguous, baselines $x^{0}\sim\mathcal{B}$ are sampled from a distribution $\mathcal{B}$ called the background.
@@ 2.3. Example-based Explanations
Intrinsic methods for deep learning. In deep learning, intrinsic interpretability can be provided by prototype-based or distance-based methods (Montenegro et al., 2022). For instance, the Explainable Deep Neural Network (xDNN) (Angelov and Soares, 2020b) and Deep Machine Reasoning (DMR) (Angelov and Soares, 2020a) define prototypes as dense data points and classify observations based on the closest prototype. The Prototype Classifier method learns representative prototypes from training data, using an autoencoder for feature extraction and classification based on latent representations (Li et al., 2018).The Prototypical Part Network (ProtoPNet) represents image parts in clusters in a latent space, which are used to predict and explain classifications (Chen et al., 2019). Additionally, the Deep k-Nearest Neighbors (DkNN) calculates neighbors at each model layer to ensure consistent predictions, offering explanations based on similar examples across the model’s entirety (Papernot and McDaniel, 2018).
@@ 2.4. Counterfactual Explanations
Counterfactual explanations (aka algorithmic recourse) provide insights into how slight changes to input features could lead to different model outcomes, aiding in tasks like model debugging and ensuring regulatory compliance (Goethals et al., 2023; Kuppa and Le-Khac, 2021). The study in (Kuppa and Le-Khac, 2021) gives an illustration of counterfactual and other four sample categories (i. e., adversarial examples, local robustness (Zhang et al., 2024), invariant samples, and uncertainty samples) through the boundaries between human analyst and a learnt model (see Fig. 4). The application of counterfactual explanations varies with the model’s complexity and includes considerations such as model transparency, type compatibility, and adherence to constraints like feasibility and causality (Wachter et al., 2017; Dodge et al., 2019; Binns et al., 2018). The concept overlaps with other areas of research such as algorithmic recourse, inverse classification, and contrastive explanations (Karimi et al., 2021; Ustun et al., 2019; Laugel et al., 2017; Dhurandhar et al., 2018).
Figure 4. Decision boundaries between human analyst and a learnt model.
@@
Diverse counterfactuals. Recent works study the generation of multiple alternative counterfactuals per input, offering a spectrum of potential changes rather than just one nearest option (Mothilal et al., 2020). This approach empowers users by offering them various ways they could potentially modify their data to achieve a preferred result (Thang et al., 2015; Nguyen et al., 2015a; Zhao et al., 2021a).
@@ 3.1. Membership Inference Attacks (MIA)
Threat model. The adversary is able to submit $x$ to the black-box model (Liu et al., 2022d; Li et al., 2022; Carlini et al., 2022; Ye et al., 2022) to receive the prediction $f(x)$ and any corresponding explanations, despite not having direct access to the model’s internals (Quan et al., 2022) (see Fig. 5). However, they are assumed to know the model’s architecture and possess an auxiliary dataset similar to the model’s training data, reflected in much of the current research on the topic (Liu et al., 2024d).
@@
Figure 6. Model-based membership inference attacks proposed in (Liu et al., 2024d).
Liu et al. (Liu et al., 2024d) proposes a model-based attack that involves four main stages: training a shadow model, extracting attribution features, training an attack model, and inferring membership (see Fig. 6). The adversary starts by training a shadow model using an auxiliary dataset that is similar to the training data of the target model. Then, attribution maps are generated for a given sample, and perturbations are applied based on these maps to observe changes in predictions. Next, the adversary trains an attack model, typically a Multi-Layer Perceptron (MLP), using the attribution features combined with other data such as loss values and one-hot encoded class information to construct features indicative of membership.
@@ 3.3. Reconstruction Attacks
  • Threat model: We consider a machine learning model $f_{t}$ that processes confidential data $x$ from a set $X_{p}$ (for instance, facial images). It employs these private inputs to generate a prediction $\hat{y}_{t}$ (such as identifying emotions). An issue arises when an attacker gains access to the target prediction $\hat{y}_{t}$ and the explanation $\phi_{t}$ (due to reasons like a data breach, interception during transmission, or sharing on social media). One scenario is to assume that the attacker only has the compromised data, an independent dataset $x\in X_{a}$, and the ability to interact with the target model via black-box (Zhao et al., 2021b). The attacker does not require additional privileged information, such as blurred versions of the images. The objective of the attacker is to develop their own inversion model $f_{a}$ to reconstruct the original image $x$ from the model’s outputs $\hat{y}_{t},\phi_{t}$). Such a reconstruction would allow them to predict sensitive information from the reconstructed image $\hat{x}_{r}$, including the possibility of re-identifying the individual from the facial emotion recognition system (Hu et al., 2022a).
@@ 3.5. Model Extraction Attacks
Research on model extraction attacks targeting explainable AI systems is emerging (Mi et al., 2024). Milli et al. (Milli et al., 2019) develops a method that leverages the discrepancy in gradient-based explanations between an original AI model and its clone, demonstrating enhanced attack efficiency. Additionally, Ulrich et al. (Aïvodji et al., 2020) designs an attack utilising counterfactual explanations to train a cloned model with greater effectiveness. Miura et al. (Miura et al., 2021) designs a data-free attack that does not require surrogate datasets in advance.
@@ 4.1. Privacy Leaks in Counterfactual Explanations
Diverse counterfactuals equip users with a range of actionable insights to potentially alter their outcomes favorably (Mothilal et al., 2020; Nguyen et al., 2023a). However, this also increases privacy risks as it may give away additional details that could be exploited for more potent attacks (Aïvodji et al., 2020). Artelt et al. (Artelt et al., 2021) identifies a key problem with counterfactual explanations: their instability to minor input variations can lead to significantly different outcomes for similar cases. Addressing this, the authors propose studying the robustness of counterfactual explanations and suggest using plausible rather than closest counterfactuals to enhance stability (Artelt and Hammer, 2020).
@@ 4.2. Causes of Membership Inference Attacks
Repeated interaction. Kumari et al. (Kumari et al., 2024) focus on repeated interactions. The author introduce attacks using explanation variance to infer data membership, modeled through a continuous-time stochastic signaling game. The study proves an optimal attack threshold exists, analyzes equilibrium conditions, and uses simulations to assess attack effectiveness in dynamic settings.
@@ 4.4. Causes of Property Inference Attacks
Regularisation techniques like dropout and ensemble learning have been shown to prevent models from memorizing private inputs, potentially reducing the risk of information leakage (Luo et al., 2021; Melis et al., 2019; Liu et al., 2022a). Despite previous findings, Luo et al. (Luo et al., 2022) reveals that incorporating dropout in neural networks at varying rates (0.2, 0.5, 0.8) actually enhances the accuracy of certain attacks. This counterintuitive result is attributed to dropout preventing overfitting by smoothing the decision boundaries, which inadvertently benefits the attack. Nevertheless, a very high dropout rate (0.8) does decrease the success rates of one attack due to underfitting and increased randomness in the model, which disrupts the linearity between inputs and outputs.
@@ 5.1. Defences with Differential Privacy
Differential privacy (DP) is a solid, mathematically based privacy standard that defines privacy loss using a quantifiable metric (Liu et al., 2024c). It does so through mechanisms that guarantee the aggregated data output will obscure the involvement of any individual record in the dataset, as established by Dwork et al. (Dwork et al., 2014). Differential privacy is usually formalized as follows (Huang et al., 2023). A randomized mechanism $M$ with domain $D$ and range $R$ achieves $\varepsilon$-differential privacy ($\varepsilon$-DP) if, for all adjacent datasets $d,d^{\prime}$ differing by one row, and for any output set $S\subseteq R$, the following inequality holds:
@@ 5.1.3. DP-Locally Linear Maps
To create differentially private Locally Linear Maps (LLM), Harder et al. (Harder et al., 2020) employs the moments accountant technique combined with differentially private stochastic gradient descent (DP-SGD) (Abadi et al., 2016). The perturbation process involves two main steps per iteration for each minibatch of size $L$: (1) Clipping the norm of the datapoint-wise gradient $h_{t}(x_{n})$ using a threshold $C$ and adding Gaussian noise to it, resulting in $\hat{h}_{t}$: $\hat{h}_{t}\leftarrow\frac{1}{L}\sum_{n=1}^{L}h_{t}(x_{n})+\mathcal{N}(0,\sigma^{2}C^{2}I)$. (2) Updating the LLM parameters in the descending direction: $W_{t+1}\leftarrow W_{t}-\eta\hat{h}_{t}$. This process ensures that the final LLM is $(\epsilon,\delta)$-differentially private. To improve the privacy-accuracy trade-off, especially for high-dimensional inputs like images, the author suggest reducing the dimensionality of the parameters by first projecting them onto a lower-dimensional space using a shared matrix $R_{m}$, and then perturbing the gradients of the projected parameters (Xue et al., 2024).